# Cross-correlation scheme for quantum optical coherence tomography based on Michelson interferometer

### v1

@@ Appendix B SPDC source
Appendix C Reconstruction of the sample structure
Following Ref. [17], a QOCT interferogram comprises three types of terms (dips or peaks) with amplitudes $V$ (we define $V>0$ for dips and $V<0$ for peaks):
Table 1: Experimental amplitudes for the terms in Fig. 5h. and their approximation by Eq. (24).
No. Position Theory Exp. value (cps) Fit (cps)
$A_{1}$ $x_{1}$ $V_{1}$ $196\pm 2$ 196
$A_{2}$ $x_{1}+{d}/{2}$ $V_{12}$ $237\pm 2$ 237
$A_{3}$ $x_{1}+d$ $V_{2}+V_{13}$ $50\pm 2$ 50
$A_{4}$ $x_{1}+{3d}/{2}$ $V_{23}+V_{14}$ $-38\pm 2$ $-40$
$A_{5}$ $x_{1}+2d$ $V_{3}+V_{24}+V_{15}$ $20\pm 2$ 17
@@ Appendix C Reconstruction of the sample structure
    $x_{1}=0\leavevmode\nobreak\ \upmu\mathrm{m},\quad x_{2}=x_{1}+d,$
@@
    $V_{kl}=A\,\mathrm{Re}\bigl{[}\tilde{r}_{k}\,e^{i\omega_{0}kd/c}\,\tilde{r}_{l}^{*}\,e^{-i\omega_{0}ld/c}\bigr{]}.$ (23)
@@
The interferogram of Fig. 5h was first fitted with a following function:
$f_{\text{fit}}(x)=C-\sum\limits_{j=1}^{5}A_{j}\exp\left[\frac{(x-x_{1}-jd/2)^{2}}{2\sigma^{2}}\right]+kx,$ (24)
with the fitting parameters: $A_{1},\dots,A_{5}$, $\sigma$, $x_{1}$, $k$ and the desired separation $d$.
The amplitudes $A_{1},\dots,A_{5}$ were then approximated by the theoretical expressions in Table 1, with free parameters $A$, $r_{1}$, $r_{2}$, $t_{1}$, and $\phi_{0}\equiv\omega_{0}d/c$. The approximation shows a good agreement with the experimental data. Finally, fixing the amplitudes to their fitted values, we refitted the interferogram terms to extract the interface separation,
$d=110.67\pm 0.11\leavevmode\nobreak\ \upmu\mathrm{m}.$

### v2

@@ Appendix B SPDC source
Appendix C Dispersion Broadening
The effect of dispersion on OCT and QOCT resolution was analyzed in Ref. [katamadze_broadband_2025]. From Eqs. (16b,c) and (21b,c) of its Supplemental Material, the FWHM of the HOM term $M_{0}(\tau)$, which determines the QOCT axial resolution, increases by a factor
$\alpha_{0}\;=\;\sqrt{1+\delta^{2}\,\Delta^{2}\,\kappa^{2}}\,$,
while the FWHM of the single‐photon Michelson term $M_{1}(\tau)$, which corresponds to the classical OCT resolution, increases by
$\alpha_{1}\;=\;\sqrt{1+\frac{\Delta_{+}^{4}\,\kappa^{2}}{4}}\,.$
Here
$\kappa\;=\;\frac{l}{c}\,\left.\frac{\partial n}{\partial\omega}\right|_{\omega_{p}/2},$
with $n(\omega)$ the refractive index of the dispersive medium and $l$ the total round‐trip path length through that medium.
Using the Sellmeier equation for silicon from Ref. [tatian_fitting_1984], one finds
$\left.\frac{\partial n}{\partial\omega}\right|_{\omega_{p}/2}\approx-1.3\times 10^{-16}\,\mathrm{s}.$
For our experimental parameters – a pump linewidth of 0.01 nm ($\delta=2\pi\times 6.9$ GHz), an SPDC bandwidth of 50 nm ($\Delta=2\pi\times 8.7$ THz), and a silicon plate thickness $l/2=5$ mm – this yields
$\alpha_{0}\approx 1+5.5\times 10^{-5}\approx 1,$ (26a)
$\alpha_{1}\approx\sqrt{1+\frac{\bigl(\delta^{2}+\Delta^{2}\bigr)^{2}\kappa^{2}}{4}}\approx 6.64.$ (26b)
Thus, a 5 mm silicon window negligibly affects the QOCT resolution but broadens the classical OCT point‐spread function by more than a factor of six.
Appendix D Data processing details
D.1 QOCT spectra fit
Since in general the QOCT interferogram spectra contain two parasitic terms, $\tilde{M}_{1}$ and $\tilde{M}_{2}$, centered at $\omega_{\text{p}}/2$ and $\omega_{\text{p}}$ respectively, we approximate $\tilde{M}_{\mathrm{MI}}^{\mathrm{(cross)}}$ and $\tilde{M}_{\mathrm{MI}}^{\mathrm{(auto)}}$ (Fig. 5d) by the sum of Gaussian functions:
$f_{\text{fit}}(\omega)=c+a_{1}\exp\left[-\frac{(\omega-\omega_{\text{p}}/2)^{2}}{2\sigma_{1}^{2}}\right]+a_{2}\exp\left[-\frac{(\omega-\omega_{\text{p}})^{2}}{2\sigma_{2}^{2}}\right].$ (27)
Because the single-count spectrum and the part of $\tilde{M}_{\mathrm{MI}}^{\mathrm{(auto)}}$ around $\omega_{\text{p}}/2$ are both affected by the same phase fluctuations and therefore coincide (Fig. 6), we fixed the standard deviation of the first Gaussian to $\sigma_{1}=0.16\,\omega_{\text{p}}$, as extracted from the Gaussian fit of the single-count spectrum (Fig. 5c, dark-purple dashed curve). Table 1 summarizes the extracted values of all other parameters with fitting errors.
Table 1: Parameters of Gaussian fit for $\tilde{M}_{\mathrm{MI}}^{\mathrm{(cross)}}$ and $\tilde{M}_{\mathrm{MI}}^{\mathrm{(auto)}}$
Parameter $\tilde{M}_{\mathrm{MI}}^{\mathrm{(cross)}}$ $\tilde{M}_{\mathrm{MI}}^{\mathrm{(auto)}}$
$c$ (kcps) 4.6 $\pm$ 1.8 4.6 $\pm$1.3
[…]
Figure 6: Overlap of interferogram spectra from Fig. 5(c, d): FFT magnitude of the QOCT coincidence-rate autocorrelation interferogram $\tilde{M}_{\mathrm{MI}}^{\mathrm{(auto)}}$ (solid red) and FFT magnitude of the OCT single-count interferogram $\tilde{M}_{\mathrm{OCT}}$ (dashed purple).
The amplitude $a_{1}$ extracted for $\tilde{M}_{\mathrm{MI}}^{\mathrm{(cross)}}$ is negligibly small ($10^{-13}\pm 10$ kcps), confirming the absence of the $\tilde{M}_{1}$ peak. The fitted values of $\sigma_{2}$ for the $\tilde{M}_{2}$ peaks in both $\tilde{M}_{\mathrm{MI}}^{\mathrm{(auto)}}$ and $\tilde{M}_{\mathrm{MI}}^{\mathrm{(cross)}}$ are $0.40\,\omega_{\text{p}}$, nearly twice as large as $\sigma_{1}$. This agrees with the expectation that the same optical path fluctuation amplitude induces twice the phase modulation depth at twice the frequency, leading to proportionally larger spectral broadening.
D.2 OCT and QOCT data processing for defectoscopy experiment
During the defectoscopy experiment, we repeated the scan eight times. Because the source intensity was not perfectly stable within and between runs, we introduced a normalization procedure to eliminate slow drifts in brightness and make the different traces comparable. Each single-count and coincidence trace was normalized by
$R_{\text{norm}}=\frac{R_{1}+\kappa_{2}R_{2}+\kappa_{3}R_{3}}{\left<R_{1}+\kappa_{2}R_{2}+\kappa_{3}R_{3}\right>},$ (28)
where $R_{1}$, $R_{2}$, and $R_{3}$ are the single-count rates from detectors D1, D2, and D3, respectively; $\langle\dots\rangle$ denotes averaging over the entire trace; and the coefficients $\kappa_{2}$, $\kappa_{3}$ were chosen such that $\langle R_{1}\rangle=2\kappa_{2}\langle R_{2}\rangle=2\kappa_{3}\langle R_{3}\rangle$.
[…]
The cross-correlation interferogram $\tilde{M}_{\mathrm{MI}}^{\mathrm{(cross)}}$ (light-blue curve in Fig. 5h) was obtained by averaging the coincidence data across all eight measurements.
D.3 Reconstruction of the sample structure
Following Ref. [li-gomezQuantumEnhancedProbing2023], a QOCT interferogram comprises three types of terms (dips or peaks) with amplitudes $V$ (we define $V>0$ for dips and $V<0$ for peaks):
@@ D.3 Reconstruction of the sample structure
    $x_{1}=0\penalty 10000\ \upmu\mathrm{m},\quad x_{2}=x_{1}+d,$
@@
    $V_{kl}=A\,\mathrm{Re}\bigl[\tilde{r}_{k}\,\operatorname{e}^{\operatorname{i}\omega_{0}kd/c}\,\tilde{r}_{l}^{*}\,\operatorname{e}^{-\operatorname{i}\omega_{0}ld/c}\bigr].$ (32)
@@
The interferogram of Fig. 5h was first fitted with the following function:
$f_{\text{fit}}(x)=C-\sum\limits_{j=1}^{5}A_{j}\exp\left[-\frac{(x-x_{1}-jd/2)^{2}}{2\sigma^{2}}\right]+kx,$ (33)
with the fitting parameters: $A_{1},\dots,A_{5}$ (enumeration corresponds to the term numbers in Fig. 5h), $\sigma$, $x_{1}$, $k$ and the desired separation $d$.
Table 2: Experimental amplitudes of the terms in Fig. 5h, and their theoretical approximation.
Theoretical Exper. Approx.
Amplitude Position approximation value value
(cps) (cps)
$A_{1}$ $x_{1}$ $V_{1}$ $196\pm 2$ 196
$A_{2}$ $x_{1}+{d}/{2}$ $V_{12}$ $237\pm 2$ 237
$A_{3}$ $x_{1}+d$ $V_{2}+V_{13}$ $50\pm 2$ 50
$A_{4}$ $x_{1}+{3d}/{2}$ $V_{23}+V_{14}$ $38\pm 2$ 40-40
$A_{5}$ $x_{1}+2d$ $V_{3}+V_{24}+V_{15}$ $20\pm 2$ 17
The amplitudes $A_{1},\dots,A_{5}$ were then approximated by the theoretical expressions in Table 2, where the free parameters $A$, $r_{1}$, $r_{2}$, $t_{1}$, and $\phi_{0}\equiv\omega_{0}d/c$ were adjusted so as to minimize the discrepancy between experiment and theory.
The approximation shows a good agreement with the experimental data. Finally, fixing the amplitudes to their approximated values, we refitted the interferogram terms to extract the interface separation, $d=110.67\pm 0.11\penalty 10000\ \upmu\mathrm{m}$.